OpenStudy (anonymous):

The area formula for a triangle is shown below. A = bh Which of the following shows this equation correctly solved for b?

OpenStudy (anonymous):

where are the options ?

OpenStudy (anonymous):

A = 1/2*base*height IF right angled

OpenStudy (anonymous):

A = bh (for b) A/h = b

OpenStudy (anonymous):

thats not correct......lamgmg90 is correct... The formula for area of a triangle is : A = 1/2bh A/(1/2)(h) = b